OpenStudy (marcelie):

help please !!! find the equation for the circle that satisfies the conditions 5) center at the origin passes through (4,7)

OpenStudy (baru):

do you know the distance formula?

OpenStudy (marcelie):

no

OpenStudy (dayakar):

distance between origin (0,0) and (x,y)=(4,7) \[\sqrt{x^{2}+y ^{2}}\]

OpenStudy (dayakar):

do u find radius = distance between origin and the point

OpenStudy (marcelie):

is midpoint formula needed here

OpenStudy (lochana):

@marcelie do you need a help?

OpenStudy (marcelie):

yes

OpenStudy (lochana):



OpenStudy (marcelie):

okay. then

OpenStudy (lochana):

now first find radius. radius^2 = (4-0)^2 + (7-0)^2= 16+49 = 65 \[radius = 65^{1/2}\]

OpenStudy (lochana):

now you have to know , there is a theorem to find formula of a circle with radius

OpenStudy (lochana):

it is x^2 + y^2 = radius^2

OpenStudy (lochana):

so the answer is x^2 + y^2 = 65

OpenStudy (marcelie):

is midpoint formula needed ?

OpenStudy (lochana):

no. you don't need to find midpoint for that. you need to know radius. that's it. wait , I will send you references

OpenStudy (lochana):

look at the diagram at left side on that page.

OpenStudy (lochana):

hope that helps. bye

OpenStudy (marcelie):

wait..

OpenStudy (marcelie):

how come on number five here they used midpt o.o

OpenStudy (lochana):

it is difference. nothing has to do with this problem

OpenStudy (marcelie):

ah ok

OpenStudy (lochana):

look closely, on number 5, they don't say center is (0,0)

OpenStudy (lochana):

(0,0) is on the curve line of circle

OpenStudy (lochana):

but in our problem, (0,0) is the center. get it?

OpenStudy (lochana):

bye

OpenStudy (dayakar):

the problem u asked is different from the example one

OpenStudy (dayakar):

@lochana ur solution is right

OpenStudy (lochana):

yes. he confused center and origin. that's why he had that problem i guess